OpenStudy (girlover):

Fan and medal!!! If log(x+3)-log(x-5)=log2,what is the value of x? A) 2 B) 13 C) 5 D) -30

OpenStudy (girlover):

@fateal @Jesstho.-.

jimthompson5910 (jim_thompson5910):

\[\Large \log(x+3)-\log(x-5) = \log(2)\] \[\Large \log\left(\frac{x+3}{x-5}\right) = \log(2)\] \[\Large \frac{x+3}{x-5} = 2\] \[\Large x+3 = 2(x-5)\] I'll let you finish up

jimthompson5910 (jim_thompson5910):

In step 2, I used rule #2 from this page http://www.purplemath.com/modules/logrules.htm

OpenStudy (triciaal):

when you subtract logs you divide the numbers you have an equation